OpenStudy (anonymous):

For the line segment whose endpoints are L (0, 1) and M (2, 8), find the y value for the point located 3 over 5 the distance from L to M.

OpenStudy (anonymous):

@ghass1978 @ganeshie8 @rational @DawnR @amistre64 @Awesome781 @BIGGHONCHO @jenniferjuice @kirbykirby @xartaan @hartnn @Matt0918 @loganmast @eliassaab @WinryRockbell @Roc07or

OpenStudy (amistre64):

what is 3/5 the distance from the xs? and from the ys?

OpenStudy (amistre64):

the xs go from 0 to 2, for a length of 2 2/5 is 1 out of 5 parts, we want 3 out of 5 parts 3*2/5 = 6/5. 0 + 6/5 = 6/5 can you work the y parts?

OpenStudy (amistre64):

the ys go from __ to __, for a length of __ __/5 is 1 out of 5 parts, we want 3 out of 5 parts 3*__/5 = __/5. 1 + __/5 = _____

OpenStudy (anonymous):

why would you put me in here :P i've never even talked in math until now

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

:P lol it's ok

OpenStudy (anonymous):

For the line segment whose endpoints are L (0, 1) and M (2, 8), find the y value for the point located 3 over 5 the distance from L to M. 5.2 3.5 4.8 1.6

OpenStudy (amistre64):

i just demonstrated a process that you can easily follow ...

OpenStudy (anonymous):

please help

OpenStudy (dawnr):

i was surprised too.. :D

OpenStudy (anonymous):

can you help me please?

OpenStudy (dawnr):

i'll show you an example and you can just follow it

OpenStudy (dawnr):

Point A is located at (0, 4) and point B is located at (−2, −3). Find the x value for the point that is 1 over 4 the distance from point A to point B. and this is how it's done: http://oi59.tinypic.com/24d4v7l.jpg

OpenStudy (dawnr):

in your case the distance it 5 and it's 3:5 instead of 1:3 everything else is pretty much the same

OpenStudy (dawnr):

is*